# Jet definition and TMD factorisation in SIDIS

### v1

@@ Abstract
Using the colour dipole picture of Deep Inelastic Scattering (DIS) and the Colour Glass Condensate effective theory, we study semi-inclusive jet production in DIS at small $x$ in the limit where the photon virtuality $Q^{2}$ is much larger than the transverse momentum squared $P_{\perp}^{2}$ of the produced jet. In this limit, the cross-section is dominated by aligned jet configurations, that is, quark-antiquark pairs in which one of the fermions — the would-be struck quark in the Breit frame — carries most of the longitudinal momentum of the virtual photon. We show that physically meaningful jet definitions in DIS are such that the effective axis of the jet sourced by the struck quark is controlled by its virtuality rather than by its transverse momentum. For such jet definitions, we show that the next-to-leading order (NLO) cross-section admits factorisation in terms of the (sea) quark transverse momentum dependent (TMD) distribution, which in turn satisfies a universal DGLAP and Sudakov evolution.
@@ Introduction.
The semi-inclusive production of a hadron or a jet in Deep Inelastic Scattering (SIDIS) is a process of fundamental importance for the study of the partonic content inside a proton or a large nucleus. It allows for an unambiguous extraction of the quark and gluon distribution functions with increasing accuracy in perturbative QCD Altarelli et al. (1979); Baier and Fey (1979); de Florian et al. (1998); Goyal et al. (2024); Bonino et al. (2024). In the small transverse momentum region of the measured hadron or jet, it is sensitive to the transverse momentum dependent (TMD) quark distribution  Collins (1993); Ji et al. (2005), whose precise extraction for both unpolarised and polarised target is one of the main goal of the Electron-Ion Collider (EIC) physics program Accardi et al. (2016); Abdul Khalek et al. (2022); Achenbach et al. (2024). In this Letter, we consider the SIDIS process in the case of a jet measurement at small $x_{\rm Bj}$, where the CGC effective theory Iancu et al. (2002); Iancu and Venugopalan (2003); Gelis et al. (2010); Kovchegov and Levin (2013) applies, and demonstrate that the factorisation of the NLO cross-section in terms of the (sea) quark TMD necessitates the use a new jet reconstruction algorithm which captures the non-trivial dynamics of the quark struck by the virtual photon.
Compared to hadron production in SIDIS, jet process has a significant advantage because it can directly probe the TMD quark distribution without involving a TMD fragmentation function. However, a precise measurement of low transverse momentum jets is very challenging, in particular because the usual $k_{t}$-type algorithms may not be applicable. Alternative approaches and detailed phenomenology applications have been carried out in the last few years Gutierrez-Reyes et al. (2018, 2019a, 2019b); Liu et al. (2019, 2020); Aschenauer et al. (2020); Page et al. (2020); Arratia et al. (2020a, 2021, b); Kang et al. (2021a, 2020, b); Andreev et al. (2022); Lai et al. (2022); Lee et al. (2023); Arratia et al. (2023); Fang et al. (2024). Together with these developments, the novel jet algorithm proposed in this paper will help to improve the precision of future jet measurements in SIDIS at the EIC and provide important probe to the nucleon/nucleus tomography in terms of the TMD quark distributions at various $x$ ranges.
@@
We work in the dipole frame (related to the Breit frame by a longitudinal boost) where the virtual photon has four-momentum $q^{\mu}=(q^{+},q^{-}=-Q^{2}/(2q^{+}),\bm{0}_{\perp})$ and a nucleon from the nucleus target has four momentum $P^{\mu}_{N}=(0,P^{-}_{N},\bm{0}_{\perp})$ in light-cone coordinates. The standard DIS variables are defined by $Q^{2}=-q^{2}$ and $x_{\rm Bj}=Q^{2}/\hat{s}$ with $\hat{s}=2q^{+}P^{-}_{N}$. At LO in the colour dipole picture of DIS at small $x_{\rm Bj}$ Kopeliovich et al. (1981); Bertsch et al. (1981); Mueller (1990); Nikolaev and Zakharov (1991), the transversely polarised virtual photon $\gamma_{\textrm{T}}^{*}$ splits into a quark-antiquark pair which interacts with the nucleus target. We denote by $k_{1,2}^{\mu}$ the four-momenta of the quark and the antiquark and define $z_{i}=k_{i}^{+}/q^{+}$. Each outgoing parton can then form a jet for which one only measures the transverse momentum $\bm{P}_{\perp}$. We are primarily interested in the limit $Q^{2}\gg P_{\perp}^{2}$. The leading power in the $P_{\perp}^{2}/Q^{2}$ expansion of the cross-section comes from aligned-jet configurations Mueller (1999); Marquet et al. (2009); Iancu et al. (2021), that is very asymmetric $q\bar{q}$ pairs such that the quark has $1-z_{1}\sim P_{\perp}^{2}/Q^{2}\ll 1$ and carries most of the longitudinal momentum of the virtual photon while the antiquark has $z_{2}\sim P_{\perp}^{2}/Q^{2}\ll 1$, or vice-versa (cf. Fig. 1). The struck fermion in the target picture is naturally the one with $z_{i}\sim 1$, yet the tagged jet with transverse momentum $\bm{P}_{\perp}$ can be either the fast ($z_{i}\sim 1$), or the slow ($z_{i}\ll 1$), fermion in the dipole picture. In the limit $Q^{2}\gg P_{\perp}^{2}$, the semi-inclusive single jet cross-section admits TMD factorisation Marquet et al. (2009) in terms of the sea quark TMD $x\mathcal{F}_{q}(x,\bm{P}_{\perp})$ at small $x$ McLerran and Venugopalan (1999); Venugopalan (1999); Mueller (1999):
@@
$\left.\frac{\mathrm{d}\sigma^{\gamma^{*}_{\rm T}+A\to j+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm LO}$ $=\frac{8\pi^{2}\alpha_{\rm em}e_{f}^{2}}{Q^{2}}x\mathcal{F}_{q}(x,\bm{P}_{\perp})\,,$ (1)
@@
where $x=x_{\rm Bj}$ as a consequence of minus momentum conservation. $\alpha_{\rm em}$ is the fine structure constant and $e_{f}$ is the fractional electric charge of a light quark. This expression covers the case where the quark jet is measured  [^1] A similar expression holds for the case where the antiquark jet is measured. See Cacciari et al. (2019) for a discussion of the unitarity of the SIDIS cross-section with jets., but it is inclusive in the jet longitudinal momentum fraction $z$. Explicit expression for the sea quark TMD in term of the quark-antiquark dipole $S$-matrix at small $x$ can be found in Xiao et al. (2017); Marquet et al. (2009); Tong et al. (2023) (see also the Supplemental Material for a brief review).
@@
At NLO in the strong coupling $\alpha_{s}$, the dominant radiative corrections to Eq. (1) are enhanced by double Sudakov logarithms Sudakov (1956); Dokshitzer et al. (1980); Parisi and Petronzio (1979); Collins et al. (1985) in the ratio of $Q^{2}/P_{\perp}^{2}$. To DLA, these Sudakov logarithms come from virtual emissions by the struck quark in the phase-space that is forbidden to the real gluon emissions, that is real gluons which would modify the structure of the final state Mueller et al. (2013a, b). Clearly, the real gluon emissions with transverse momenta such that $k_{g\perp}^{2}\gg P_{\perp}^{2}$ must be forbidden. On the other hand, virtual emissions with $k_{g\perp}^{2}\geq Q^{2}$ do not have any double logarithmic support (the UV logarithmic divergence cancels above the hard scale $Q^{2}$ among virtual graphs Caucal et al. (2024)). The transverse phase-space which contributes to the Sudakov double logarithm is therefore $P_{\perp}^{2}\ll k_{g\perp}^{2}\ll Q^{2}$. Concerning the longitudinal phase space, we observe that the gluon emission must have a formation time $\tau_{g}=1/k_{g}^{-}$ larger than the coherence time $\tau_{\gamma}=1/|q^{-}|$ of the virtual photon Beuf (2014); Iancu et al. (2015); Hatta and Iancu (2016); Ducloué et al. (2019). In the context of high energy factorisation, this condition is imposed on any gluon emission which does not contribute to the (collinearly improved) BK/JIMWLK evolution Balitsky (1996); Kovchegov (1999); Jalilian-Marian et al. (1997, 1998); Kovner et al. (2000); Weigert (2002); Iancu et al. (2001a, b); Ferreiro et al. (2002) of the target wave-function. In terms of $z_{g}=k_{g}^{+}/q^{+}$, this gives the constraint $z_{g}\gg k_{g\perp}^{2}/Q^{2}$. We thus end up with the following integral for the Sudakov double logarithm,
@@
Using this criterion at NLO, where the final state is made of two partons, the quark $i$ and gluon $j$, one sees that these two partons are clustered within the same jet if $\theta_{ij}\leq R\theta_{q}$ with $\theta_{q}\sim Q/q^{+}$, in agreement with the dynamics of the fast virtual quark that we have just elucidated. This is a consequence of both the normalisation of the distance measure with $Q^{2}$ and the choice of $p$. We assume the jet radius $R$ to be small, but not too small, so that an additional $\ln(1/R)$ resummation is not needed Dasgupta et al. (2015)  [^3] For jet production with a very small-$R$, one needs to introduce the so-called semi-inclusive jet function Kang et al. (2016) and the factorisation will be different from what we discuss in this paper..
For $p=1$ and in the fragmentation region of the struck quark in the Breit frame, the algorithm is akin to generalised $k_{t}$ algorithms originally designed for $e^{+}e^{-}$ annihilation Bartel et al. (1986); Bethke et al. (1988); Catani et al. (1991); Cacciari et al. (2012) and subsequently extended to DIS in Catani et al. (1992); Webber (1993) (cf Supplemental Material). However, unlike the latter, which are only defined in the Breit frame, Eq. (5) is longitudinally invariant by boost along the $\gamma^{*}$-$A$ collision axis (see also Arratia et al. (2021)), so that $d_{ij},d_{iB}$ can be computed either in the Breit frame, or the dipole frame. This makes the distance measure Eq. (5) more convenient for higher order calculations and jet studies in DIS.
@@
For the inclusive production of jets in DIS, that is, in situations where the jet transverse momenta are not measured, but only their number, it was known since Catani et al. (1992); Webber (1993) that special jet definitions are needed to guarantee the validity of the standard collinear factorisation for the jet cross-section. In what follows, we would like to demonstrate a similar property for the semi-inclusive production of a jet with a given $P_{\perp}$. Specifically, we will show that the NLO corrections to the jet cross-section (as computed in the dipole picture) are consistent with TMD factorisation — in the sense that they correctly generate the expected DGLAP Gribov and Lipatov (1972); Altarelli and Parisi (1977); Dokshitzer (1977) and CSS Collins and Soper (1981, 1982); Collins et al. (1985); Collins (2023) evolutions of the quark TMD — if and only if the quark and the gluon jets are separated from each other by using the jet distance (5) with $p=1$. To show that, we consider a relatively hard jet, with $Q^{2}\gg P_{\perp}^{2}\gg Q_{s}^{2}$. In this regime and to lowest order in $Q_{s}^{2}/P_{\perp}^{2}$, the transverse momentum of the measured jet is either given by the recoil from a hard gluon emission with $\bm{k_{g\perp}}\sim-\bm{P}_{\perp}$ (real contribution) or by the target itself (virtual contribution).
@@
×$\times\frac{2z_{1}\left[(P_{\perp}^{2}+\bar{Q}^{2})^{2}+z_{g}^{2}P_{\perp}^{4}+z_{1}^{2}\bar{Q}^{4}\right]}{P_{\perp}^{2}\left[P_{\perp}^{2}+\bar{Q}^{2}\right]^{3}}x_{q}\mathcal{F}_{q}(x_{q},\bm{\ell}_{\perp})\,,$ (7)
with $\bar{Q}^{2}=z_{1}z_{g}Q^{2}$ and $x_{q}=(M_{qg}^{2}+Q^{2})/\hat{s}$. The contribution of (7) to SIDIS is obtained by integrating out the kinematics of the unmeasured jet over the region in phase-space where the quark and the gluon form two separated jets. Using the jet definition Eq. (5) with $p=1$, the associated phase space constraint reads $P_{\perp}^{2}\geq R^{2}Q^{2}z_{1}^{2}z_{g}^{2}$. The subsequent integrals over $z_{1}$ and $z_{g}$ are controlled by $z_{1}\simeq 1$ and $z_{g}\lesssim P_{\perp}/(RQ)\ll 1$, so that only very soft gluons contribute. Such soft gluons can be transferred into the target wavefunction Iancu et al. (2022a, b); Hauksson et al. (2024); Caucal and Iancu (2024) via a change of variable $z_{g}\to\xi$, with
@@
where we have also used $M_{qg}^{2}=P_{\perp}^{2}/(z_{1}z_{g})$ and $z_{1}\simeq 1$. The integral over $\xi$ is restricted to $x_{\star}<\xi<1-RP_{\perp}/Q$, where the upper limit is introduced by the jet constraint $z_{g}\lesssim P_{\perp}/(RQ)$, while the lower limit comes from the separation of the longitudinal phase-space between the NLO impact factor and the high energy evolution. This separation is more precisely written as $z_{g}\geq x_{\star}P_{\perp}^{2}/Q^{2}$, with $x_{\star}$ an arbitrary factorisation scale which obeys $x_{\star}\ll 1$ and $\alpha_{s}\ln(1/x_{\star})\ll 1$. Finally, the integral over the momentum imbalance $\ell_{\perp}\ll P_{\perp}$ builds the quark PDF $xf_{q}$ on the resolution scale of the hard jet:
$xf_{q}(x,P_{\perp}^{2})=\int_{\Lambda^{2}}^{P_{\perp}^{2}}{\mathrm{d}^{2}\bm{\ell}_{\perp}}\,x\mathcal{F}_{q}(x,\bm{\ell}_{\perp})\,,$ (9)
@@
×$\times\frac{\alpha_{s}}{2\pi^{2}}\frac{1}{P_{\perp}^{2}}\int_{x_{\star}}^{1-\frac{RP_{\perp}}{Q}}{\mathrm{d}\xi}P_{qq}(\xi)\frac{x_{\rm Bj}}{\xi}f_{q}\left(\frac{x_{\rm Bj}}{\xi},P_{\perp}^{2}\right)\,,$ (10)
with $P_{qq}(\xi)=C_{F}(1+\xi^{2})/(1-\xi)$ the unregularised $q\to qg$ splitting function. In Eq. (10) we included a factor of two to account for the fact that the tagged jet can be generated by either the fast quark ($z_{1}\simeq 1$), or the slow gluon ($z_{g}\ll 1$). The NLO correction in Eq. (10) exhibits TMD factorisation, as anticipated: the expression in the second line can be interpreted as an evolution of the quark TMD. The singularity of $P_{qq}(\xi)$ at $\xi=1$ introduces a logarithmic sensitivity to the upper limit, that ca be isolated with the help of the plus prescription:
$\left.x\mathcal{F}_{q}^{(1)}(x,\bm{P}_{\perp},Q^{2})\right|_{R}=\frac{\alpha_{s}}{2\pi^{2}}\frac{1}{P_{\perp}^{2}}\!\int_{x_{\star}}^{1}{\mathrm{d}\xi}P_{qq}^{(+)}(\xi)\frac{x}{\xi}f_{q}\left(\frac{x}{\xi},P_{\perp}^{2}\right)$
$\quad\qquad\qquad+\frac{\alpha_{s}C_{F}}{2\pi^{2}}\frac{1}{P_{\perp}^{2}}\ln\left(\frac{Q^{2}}{R^{2}P_{\perp}^{2}}\right)xf_{q}(x,P_{\perp}^{2})\,,$ (11)
where $P_{qq}^{(+)}(\xi)$ differs from $P_{qq}(\xi)$ only via the replacement $(1-\xi)\to(1-\xi)_{+}$ in the denominator. We recognise in Eq. (11) one (real) step in the DGLAP+CSS evolution of the quark TMD. (See also Hauksson et al. (2024); Caucal and Iancu (2024) for similar arguments.) The $1/P_{\perp}^{2}$ behaviour comes from the hard $q\to qg$ splitting leading to two well separated jets. The precise coefficient of the Sudakov logarithm in the second line is a consequence of choosing $p=1$ in our jet distance measure (5). If one were using another jet definition from that class, or a definition with an effective jet axis set by the angle $\theta_{0}$ in Fig. 1, the upper limit of the $\xi$ integral in Eq. (10) would change (e.g., it would be $\xi<1-R^{2}P_{\perp}^{2}/Q^{2}$ for $p=0$), which would in turn modify the normalisation of the Sudakov logarithm. With $p=-1$ in Eq. (5), the quark and the gluon would typically be clustered within the same jet, so that the TMD evolution that we have just unveiled would not be resolved within the wide jet formed by the struck quark.

### v2

# Jet Definition and Transverse-Momentum-Dependent Factorization in Semi-Inclusive Deep-Inelastic Scattering

@@ Abstract
Using the colour dipole picture of Deep Inelastic Scattering (DIS) and the Colour Glass Condensate effective theory, we study semi-inclusive jet production in DIS at small $x$ in the limit where the photon virtuality $Q^{2}$ is much larger than the transverse momentum squared $P_{\perp}^{2}$ of the produced jet. In this limit, the cross-section is dominated by aligned jet configurations, that is, quark-antiquark pairs in which one of the fermions — the would-be struck quark in the Breit frame — carries most of the longitudinal momentum of the virtual photon. We show that physically meaningful jet definitions in DIS are such that the effective axis of the jet sourced by the struck quark is controlled by its virtuality rather than by its transverse momentum. For such jet definitions, we show that the next-to-leading order (NLO) cross-section admits factorisation in terms of the (sea) quark transverse momentum dependent (TMD) distribution, which in turn satisfies a universal Dokshitzer-Gribov-Lipatov-Altarelli-Parisi and Sudakov evolution.
@@ Introduction.
The semi-inclusive production of a hadron or a jet in Deep Inelastic Scattering (SIDIS) is a process of fundamental importance for the study of the partonic content inside a proton or a large nucleus. It allows for an unambiguous extraction of the quark and gluon distribution functions with increasing accuracy in perturbative QCD Altarelli et al. (1979); Baier and Fey (1979); de Florian et al. (1998); Abelof et al. (2016); Goyal et al. (2024); Bonino et al. (2024). In the small transverse momentum region of the measured hadron or jet, it is sensitive to the transverse momentum dependent (TMD) quark distribution  Collins (1993); Ji et al. (2005); Collins (2023); Boussarie et al. (2023), whose precise extraction for both unpolarised and polarised target is one of the main goal of the Electron-Ion Collider (EIC) physics program Accardi et al. (2016); Abdul Khalek et al. (2022); Achenbach et al. (2024). In this Letter, we consider the SIDIS process in the case of a jet measurement at small $x_{\rm Bj}$, where the Colour Glass Condensate (CGC) effective theory Iancu et al. (2002); Iancu and Venugopalan (2003); Gelis et al. (2010); Kovchegov and Levin (2013) applies, and demonstrate that the factorisation of the NLO cross-section in terms of the (sea) quark TMD necessitates the use a new jet reconstruction algorithm which captures the non-trivial dynamics of the quark struck by the virtual photon.
Compared to hadron production in SIDIS, jet production has the advantage that it can directly probe the TMD quark distribution without involving a TMD fragmentation function. However, a precise measurement of low transverse momentum jets is very challenging, in particular because the usual $k_{t}$-type algorithms may not be applicable. Alternative approaches and detailed phenomenology applications have been carried out in the last few years Gutierrez-Reyes et al. (2018, 2019a, 2019b); Liu et al. (2019, 2020); Aschenauer et al. (2020); Page et al. (2020); Arratia et al. (2020a, 2021, b); Kang et al. (2021a, 2020, b); Andreev et al. (2022); Lai et al. (2022); Lee et al. (2023); Arratia et al. (2023); Fang et al. (2024). Together with these developments, the novel jet algorithm proposed in this paper will help to improve the precision of future jet measurements in SIDIS at the EIC and provide important probe to the nucleon/nucleus tomography in terms of the TMD quark distributions at various $x$ ranges.
@@
We work in the dipole frame (related to the Breit frame by a longitudinal boost) where the virtual photon has four-momentum $q^{\mu}=(q^{+},q^{-}=-Q^{2}/(2q^{+}),\bm{0}_{\perp})$ and a nucleon from the nucleus target has four momentum $P^{\mu}_{N}=(0,P^{-}_{N},\bm{0}_{\perp})$ in light-cone coordinates. The standard DIS variables are defined by $Q^{2}=-q^{2}$ and $x_{\rm Bj}=Q^{2}/\hat{s}$ with $\hat{s}=2q^{+}P^{-}_{N}$. At LO in the colour dipole picture of DIS at small $x_{\rm Bj}$ Kopeliovich et al. (1981); Bertsch et al. (1981); Mueller (1990); Nikolaev and Zakharov (1991), the transversely polarised virtual photon $\gamma_{\textrm{T}}^{*}$ splits into a quark-antiquark pair which interacts with the nucleus target. We denote by $k_{1,2}^{\mu}$ the four-momenta of the quark and the antiquark and define $z_{i}=k_{i}^{+}/q^{+}$. Each outgoing parton can then form a jet for which one only measures the transverse momentum $\bm{P}_{\perp}$. We are primarily interested in the limit $Q^{2}\gg P_{\perp}^{2}$. The leading power in the $P_{\perp}^{2}/Q^{2}$ expansion of the cross-section comes from aligned-jet configurations Mueller (1999); Marquet et al. (2009); Iancu et al. (2021), that is very asymmetric $q\bar{q}$ pairs such that the quark has $1-z_{1}\sim P_{\perp}^{2}/Q^{2}\ll 1$ and carries most of the longitudinal momentum of the virtual photon while the antiquark has $z_{2}\sim P_{\perp}^{2}/Q^{2}\ll 1$, or vice-versa (cf. Fig. 1). The struck fermion in the target picture is naturally the one with $z_{i}\sim 1$, yet the tagged jet with transverse momentum $\bm{P}_{\perp}$ can be either the fast ($z_{i}\sim 1$), or the slow ($z_{i}\ll 1$), fermion in the dipole picture. In the Breit frame, this would give either a forward jet in the direction of the photon, or a backward jet close to the beam remnants.
In the limit $Q^{2}\gg P_{\perp}^{2}$, the semi-inclusive single jet cross-section admits TMD factorisation Marquet et al. (2009) in terms of the sea quark TMD $x\mathcal{F}_{q}(x,\bm{P}_{\perp})$ at small $x$ McLerran and Venugopalan (1999); Venugopalan (1999); Mueller (1999):
@@
$\left.\frac{\mathrm{d}\sigma^{\gamma^{*}_{\rm T}+A\to j+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm LO}$ $=\frac{8\pi^{2}\alpha_{\rm em}e_{f}^{2}}{Q^{2}}x\mathcal{F}_{q}^{(0)}(x,\bm{P}_{\perp})\,,$ (1)
@@
where $x=x_{\rm Bj}$ as a consequence of minus momentum conservation. The superscript $(0)$ for $\mathcal{F}_{q}$ refers to a LO approximation for the quark TMD including its high energy BK/JIMWLK evolution Balitsky (1996); Kovchegov (1999); Jalilian-Marian et al. (1997, 1998); Kovner et al. (2000); Weigert (2002); Iancu et al. (2001a, b); Ferreiro et al. (2002) down to the value of $x$ of interest. $\alpha_{\rm em}$ is the fine structure constant and $e_{f}$ is the fractional electric charge of a light quark. This expression covers the case where the quark jet is measured  [^1] A similar expression holds for the case where the antiquark jet is measured. See Cacciari:2019qjx for a discussion of the unitarity of the SIDIS cross-section with jets., but it is inclusive in the jet longitudinal momentum fraction $z$. Explicit expression for the sea quark TMD in term of the quark-antiquark dipole $S$-matrix at small $x$ can be found in Xiao et al. (2017); Marquet et al. (2009); Tong et al. (2023) (see also the Supplemental Material  [^2] See Supplemental Material, which includes Refs. Newman and Wing (2014); Chekanov et al. (2002, 2007); Abramowicz et al. (2010); Aktas et al. (2007); Andreev et al. (2017); Abt et al. (2023); Liu and Petriello (2013); Catani et al. (1993); Blazey et al. (2000); Ducloué et al. (2020); Dominguez et al. (2011); Bergabo and Jalilian-Marian (2022); Iancu and Mulian (2023); Bergabo and Jalilian-Marian (2023a, b); Becher and Neubert (2011); Becher et al. (2013); Echevarria et al. (2012, 2014); Chiu et al. (2012); Ebert et al. (2020); Sun and Yuan (2013), for a comparison of the various jet definitions used in DIS, details on the derivation of Eq. (12) and a general discussion on TMD factorisation for jet production in SIDIS at moderate values for $x$. for a brief review).
@@
At NLO in the strong coupling $\alpha_{s}$, the dominant radiative corrections to Eq. (1) are enhanced by double Sudakov logarithms Sudakov (1956); Dokshitzer et al. (1980); Parisi and Petronzio (1979); Collins et al. (1985) in the ratio of $Q^{2}/P_{\perp}^{2}$. To DLA, these Sudakov logarithms come from virtual emissions by the struck quark in the phase-space that is forbidden to the real gluon emissions, that is real gluons which would modify the structure of the final state Mueller et al. (2013a, b). Clearly, the real gluon emissions with transverse momenta such that $k_{g\perp}^{2}\gg P_{\perp}^{2}$ must be forbidden. On the other hand, virtual emissions with $k_{g\perp}^{2}\geq Q^{2}$ do not have any double logarithmic support (the UV logarithmic divergence cancels above the hard scale $Q^{2}$ among virtual graphs Caucal et al. (2024)). The transverse phase-space which contributes to the Sudakov double logarithm is therefore $P_{\perp}^{2}\ll k_{g\perp}^{2}\ll Q^{2}$. Concerning the longitudinal phase space, we observe that the gluon emission must have a formation time $\tau_{g}=1/k_{g}^{-}$ larger than the coherence time $\tau_{\gamma}=1/|q^{-}|$ of the virtual photon Beuf (2014); Iancu et al. (2015); Hatta and Iancu (2016); Ducloué et al. (2019). In the context of high energy factorisation, this condition is imposed on any gluon emission which does not contribute to the (collinearly improved) Balitsky-Kovchegov (BK) and Jalilian-Marian-Iancu-McLerran-Weigert-Leonidov-Kovner (JIMWLK) evolutions Balitsky (1996); Kovchegov (1999); Jalilian-Marian et al. (1997, 1998); Kovner et al. (2000); Weigert (2002); Iancu et al. (2001a, b); Ferreiro et al. (2002) of the target wave-function. In terms of $z_{g}=k_{g}^{+}/q^{+}$, this gives the constraint $z_{g}\gg k_{g\perp}^{2}/Q^{2}$. We thus end up with the following integral for the Sudakov double logarithm,
@@
Using this criterion at NLO, where the final state is made of two partons, the quark $i$ and gluon $j$, one sees that these two partons are clustered within the same jet if $\theta_{ij}\leq R\theta_{q}$ with $\theta_{q}\sim Q/q^{+}$, in agreement with the dynamics of the fast virtual quark that we have just elucidated. This is a consequence of both the normalisation of the distance measure with $Q^{2}$ and the choice of $p$. We assume the jet radius $R$ to be small, but not too small, so that an additional $\ln(1/R)$ resummation is not needed Dasgupta et al. (2015)  [^4] For jet production with a very small-$R$, one needs to introduce the so-called semi-inclusive jet function Kang:2016mcy and the factorisation will be different from what we discuss in this paper..
For $p=1$ and in the fragmentation region of the struck quark in the Breit frame, the algorithm is akin to generalised $k_{t}$ algorithms originally designed for $e^{+}e^{-}$ annihilation Bartel et al. (1986); Bethke et al. (1988); Catani et al. (1991); Cacciari et al. (2012) and subsequently extended to DIS in Catani et al. (1992); Webber (1993) (cf Supplemental Material). However, unlike the latter, which are only defined in the Breit frame, Eq. (5) is longitudinally invariant by boost along the $\gamma^{*}$-$A$ collision axis (see also Arratia et al. (2021); van Beekveld and Ferrario Ravasio (2024)), so that $d_{ij},d_{iB}$ can be computed either in the Breit frame, or the dipole frame. This makes the distance measure Eq. (5) more convenient for higher order calculations and jet studies in DIS.
@@
For the inclusive production of jets in DIS, that is, in situations where the jet transverse momenta are not measured, but only their number, it was known since Catani et al. (1992); Webber (1993) that special jet definitions are needed to guarantee the validity of the standard collinear factorisation for the jet cross-section. In what follows, we would like to demonstrate a similar property for the semi-inclusive production of a jet with a given $P_{\perp}$. Specifically, we will show that the NLO corrections to the jet cross-section (as computed in the dipole picture) are consistent with TMD factorisation — in the sense that they correctly generate the expected Dokshitzer-Gribov-Lipatov-Altarelli-Parisi (DGLAP) Gribov and Lipatov (1972); Altarelli and Parisi (1977); Dokshitzer (1977) and Collins-Soper-Sterman (CSS) Collins and Soper (1981, 1982); Collins et al. (1985); Collins (2023) evolutions of the quark TMD — if and only if the quark and the gluon jets are separated from each other by using the jet distance (5) with $p=1$. To show that, we consider a relatively hard jet, with $Q^{2}\gg P_{\perp}^{2}\gg Q_{s}^{2}$. In this regime and to lowest order in $Q_{s}^{2}/P_{\perp}^{2}$, the transverse momentum of the measured jet is either given by the recoil from a hard gluon emission with $\bm{k_{g\perp}}\sim-\bm{P}_{\perp}$ (real contribution) or by the target itself (virtual contribution).
@@
×$\times\frac{2z_{1}\left[(P_{\perp}^{2}+\bar{Q}^{2})^{2}+z_{g}^{2}P_{\perp}^{4}+z_{1}^{2}\bar{Q}^{4}\right]}{P_{\perp}^{2}\left[P_{\perp}^{2}+\bar{Q}^{2}\right]^{3}}x_{q}\mathcal{F}_{q}^{(0)}(x_{q},\bm{\ell}_{\perp})\,,$ (7)
with $\bar{Q}^{2}=z_{1}z_{g}Q^{2}$ and $x_{q}=(M_{qg}^{2}+Q^{2})/\hat{s}$. The contribution of (7) to SIDIS is obtained by integrating out the kinematics of the unmeasured jet over the region in phase-space where the quark and the gluon form two separated jets. Using the jet definition Eq. (5) with $p=1$, the associated phase space constraint reads $P_{\perp}^{2}\geq R^{2}Q^{2}z_{1}^{2}z_{g}^{2}$. The subsequent integrals over $z_{1}$ and $z_{g}$ are controlled by $z_{1}\simeq 1$ and $z_{g}\lesssim P_{\perp}/(RQ)\ll 1$, so only relatively soft gluons contribute. Such soft gluons can be transferred into the target wavefunction Iancu et al. (2022a, b); Hauksson et al. (2024); Caucal and Iancu (2024) via a change of variable $z_{g}\to\xi$, with
@@
where we have also used $M_{qg}^{2}=P_{\perp}^{2}/(z_{1}z_{g})$ and $z_{1}\simeq 1$. The integral over $\xi$ is restricted to $x_{\rm Bj}<\xi<1-RP_{\perp}/Q$, where the upper limit is introduced by the jet constraint $z_{g}\lesssim P_{\perp}/(RQ)$, while the lower limit comes from the condition $x_{q}\leq 1$. Note that Eq. (7) is not singular as $z_{g}\to 0$ meaning that the contribution from Eq. (7) to the NLO cross-section does not overlap with the high energy evolution already accounted for through the $x$ dependence of $x\mathcal{F}_{q}^{(0)}$ in Eq. (1). Finally, the integral over the momentum imbalance $\ell_{\perp}\ll P_{\perp}$ builds the quark PDF $xf_{q}$ on the resolution scale of the hard jet:
$xf_{q}^{(0)}(x,P_{\perp}^{2})=\int_{\Lambda^{2}}^{P_{\perp}^{2}}{\mathrm{d}^{2}\bm{\ell}_{\perp}}\,x\mathcal{F}_{q}^{(0)}(x,\bm{\ell}_{\perp})\,,$ (9)
@@
×$\times\frac{\alpha_{s}}{2\pi^{2}}\frac{1}{P_{\perp}^{2}}\int_{x_{\rm Bj}}^{1-\frac{RP_{\perp}}{Q}}{\mathrm{d}\xi}P_{qq}(\xi)\frac{x_{\rm Bj}}{\xi}f_{q}^{(0)}\left(\frac{x_{\rm Bj}}{\xi},P_{\perp}^{2}\right)\,,$ (10)
with $P_{qq}(\xi)=C_{F}(1+\xi^{2})/(1-\xi)$ the unregularised $q\to qg$ splitting function. In Eq. (10) we included a factor of two to account for the fact that the tagged jet can be generated by either the fast quark ($z_{1}\simeq 1$), or the slow gluon ($z_{g}\ll 1$). The NLO correction in Eq. (10) exhibits TMD factorisation, as anticipated: the expression in the second line can be interpreted as an evolution of the quark TMD. The singularity of $P_{qq}(\xi)$ at $\xi=1$ introduces a logarithmic sensitivity to the upper limit, that can be isolated with the help of the plus prescription:
$x\mathcal{F}_{q}^{(1)}(x,\bm{P}_{\perp},Q^{2})\Big{|}_{R}=\frac{\alpha_{s}}{2\pi^{2}}\frac{1}{P_{\perp}^{2}}\!\int_{x}^{1}\!{\mathrm{d}\xi}P_{qq}^{(+)}(\xi)\frac{x}{\xi}f_{q}^{(0)}\!\left(\frac{x}{\xi},P_{\perp}^{2}\right)$
$\quad\qquad\qquad+\frac{\alpha_{s}C_{F}}{2\pi^{2}}\frac{1}{P_{\perp}^{2}}\ln\left(\frac{Q^{2}}{R^{2}P_{\perp}^{2}}\right)xf_{q}^{(0)}(x,P_{\perp}^{2})\,,$ (11)
where $P_{qq}^{(+)}(\xi)$ differs from $P_{qq}(\xi)$ only via the replacement $(1-\xi)\to(1-\xi)_{+}$ in the denominator. We recognise in Eq. (11) one (real) step in the DGLAP+CSS evolution of the quark TMD. (See also Hauksson et al. (2024); Caucal and Iancu (2024) for similar arguments.) As shown in the SM, Eq. (11) is also the standard result for the one-loop contribution of the collinear gluon radiation to the quark TMD at moderate $x$ and $P_{\perp}\gg\Lambda$ Boussarie et al. (2023). Recovering this well known result is a non-trivial check of our jet definition. Indeed, the precise coefficient of the Sudakov logarithm in the second line is a consequence of choosing $p=1$ in our jet distance measure (5). If one were using another jet definition from that class, or a definition with an effective jet axis set by the angle $\theta_{0}$ in Fig. 1, the upper limit of the $\xi$ integral in Eq. (10) would change (e.g., it would be $\xi<1-R^{2}P_{\perp}^{2}/Q^{2}$ for $p=0$), which would in turn modify the normalisation of the Sudakov logarithm. With $p=-1$ in Eq. (5), the quark and the gluon would typically be clustered within the same jet, so that the TMD evolution that we have just unveiled would not be resolved within the wide jet formed by the struck quark.